# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.28 15.08 14.55
Bleu score std 14.76 16.23 15.04
Bleu 1gram precision mean 14.34 15.15 14.63
Bleu 1gram precision std 14.82 16.29 15.11
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 177.82 179.16 178.20
Bleu ref len std 65.99 65.72 66.07
Bleu pred len mean 178.21 179.61 178.54
Bleu pred len std 66.06 65.78 66.13
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 14.77 15.59 15.05
Chrf score std 15.02 16.49 15.29
Ter score mean 102.45 101.53 102.09
Ter score std 18.88 20.35 18.95
Chrfpp score mean 14.77 15.59 15.05
Chrfpp score std 15.02 16.49 15.29
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 11.47 11.83 11.57
Bleu score std 10.75 11.92 11.66
Bleu 1gram precision mean 11.50 11.86 11.61
Bleu 1gram precision std 10.78 11.94 11.69
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 183.01 183.18 183.27
Bleu ref len std 66.11 66.21 66.10
Bleu pred len mean 185.58 185.77 185.77
Bleu pred len std 66.51 66.57 66.44
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 11.95 12.32 12.05
Chrf score std 11.00 12.15 11.88
Ter score mean 106.35 106.05 106.22
Ter score std 15.12 16.63 15.96
Chrfpp score mean 11.95 12.32 12.05
Chrfpp score std 11.00 12.15 11.88
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 10.24 12.43 12.31
Bleu score std 11.68 14.39 13.71
Bleu 1gram precision mean 10.29 12.51 12.38
Bleu 1gram precision std 11.72 14.45 13.76
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 75.92 80.18 79.39
Bleu ref len std 31.48 32.63 32.60
Bleu pred len mean 79.72 84.07 83.38
Bleu pred len std 34.40 35.19 35.41
Bleu bp mean 0.98 0.99 0.99
Bleu bp std 0.11 0.08 0.09
Bleu ratio mean 1.04 1.05 1.05
Bleu ratio std 0.14 0.12 0.13
Chrf score mean 10.82 13.08 12.98
Chrf score std 12.05 14.82 14.16
Ter score mean 101.05 98.60 98.46
Ter score std 19.34 21.01 19.97
Chrfpp score mean 10.82 13.08 12.98
Chrfpp score std 12.05 14.82 14.16
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 13.67 14.77 14.40
Bleu score std 15.29 15.79 16.04
Bleu 1gram precision mean 13.75 14.86 14.48
Bleu 1gram precision std 15.31 15.84 16.08
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 75.67 81.10 78.32
Bleu ref len std 34.88 34.04 33.85
Bleu pred len mean 78.80 84.24 81.61
Bleu pred len std 36.63 35.88 35.88
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.05 0.04 0.04
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.08 0.07 0.07
Chrf score mean 14.34 15.42 15.06
Chrf score std 15.65 16.10 16.38
Ter score mean 97.26 95.50 96.19
Ter score std 22.32 21.46 22.40
Chrfpp score mean 14.34 15.42 15.06
Chrfpp score std 15.65 16.10 16.38
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.21 18.44 18.60
Bleu score std 18.76 19.49 19.25
Bleu 1gram precision mean 17.35 18.54 18.70
Bleu 1gram precision std 18.95 19.62 19.38
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 83.21 85.58 85.50
Bleu ref len std 36.03 35.24 35.33
Bleu pred len mean 86.68 88.99 88.96
Bleu pred len std 37.81 37.00 37.10
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06
Chrf score mean 17.93 19.16 19.32
Chrf score std 19.10 19.82 19.56
Ter score mean 93.01 90.78 90.55
Ter score std 24.89 24.17 24.00
Chrfpp score mean 17.93 19.16 19.32
Chrfpp score std 19.10 19.82 19.56
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.50 14.90 14.70
Bleu score std 15.36 16.07 15.52
Bleu 1gram precision mean 14.58 14.98 14.78
Bleu 1gram precision std 15.49 16.20 15.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 84.04 84.57 84.06
Bleu ref len std 35.54 35.28 35.58
Bleu pred len mean 87.66 88.19 87.69
Bleu pred len std 37.74 37.49 37.81
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06
Chrf score mean 15.12 15.52 15.32
Chrf score std 15.65 16.35 15.81
Ter score mean 95.18 94.49 94.76
Ter score std 20.33 20.54 20.07
Chrfpp score mean 15.12 15.52 15.32
Chrfpp score std 15.65 16.35 15.81
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 12.04 12.98 12.21
Bleu score std 13.46 14.42 13.42
Bleu 1gram precision mean 12.08 13.03 12.24
Bleu 1gram precision std 13.51 14.47 13.46
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 86.46 86.23 86.48
Bleu ref len std 36.21 36.19 36.24
Bleu pred len mean 92.34 91.99 92.37
Bleu pred len std 39.90 39.68 39.94
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.02 0.01
Bleu ratio mean 1.07 1.07 1.07
Bleu ratio std 0.08 0.09 0.08
Chrf score mean 12.76 13.74 12.94
Chrf score std 13.86 14.81 13.83
Ter score mean 99.34 98.29 99.13
Ter score std 20.49 21.52 20.54
Chrfpp score mean 12.76 13.74 12.94
Chrfpp score std 13.86 14.81 13.83
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 22.68 25.41 23.57
Bleu score std 20.79 23.06 21.92
Bleu 1gram precision mean 22.75 25.47 23.64
Bleu 1gram precision std 20.83 23.10 21.96
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 203.28 234.13 207.02
Bleu ref len std 84.53 89.52 84.94
Bleu pred len mean 204.82 235.82 208.57
Bleu pred len std 85.31 90.00 85.73
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 23.33 26.10 24.25
Chrf score std 21.07 23.29 22.15
Ter score mean 91.40 88.33 90.51
Ter score std 25.52 27.99 26.91
Chrfpp score mean 23.33 26.10 24.25
Chrfpp score std 21.07 23.29 22.15
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.34 29.55 25.83
Bleu score std 20.94 24.54 21.80
Bleu 1gram precision mean 24.42 29.60 25.90
Bleu 1gram precision std 20.98 24.57 21.84
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 237.17 246.51 243.30
Bleu ref len std 90.05 87.83 87.19
Bleu pred len mean 238.76 248.14 244.78
Bleu pred len std 91.05 88.54 87.84
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.01 1.01
Bleu ratio std 0.03 0.02 0.03
Chrf score mean 25.05 30.30 26.55
Chrf score std 21.09 24.71 22.03
Ter score mean 89.75 83.28 87.62
Ter score std 26.10 29.48 26.32
Chrfpp score mean 25.05 30.30 26.55
Chrfpp score std 21.09 24.71 22.03
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.45 28.83 25.73
Bleu score std 22.85 26.47 23.16
Bleu 1gram precision mean 24.51 28.89 25.79
Bleu 1gram precision std 22.88 26.50 23.19
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 245.51 252.04 251.56
Bleu ref len std 96.46 91.08 91.27
Bleu pred len mean 245.92 252.59 251.95
Bleu pred len std 96.66 91.19 91.33
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 25.04 29.44 26.35
Chrf score std 22.96 26.59 23.33
Ter score mean 89.16 83.80 87.36
Ter score std 27.90 31.46 27.52
Chrfpp score mean 25.04 29.44 26.35
Chrfpp score std 22.96 26.59 23.33
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.49 26.86 24.57
Bleu score std 21.92 23.53 22.03
Bleu 1gram precision mean 24.58 26.93 24.65
Bleu 1gram precision std 21.96 23.57 22.07
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 246.19 249.56 246.63
Bleu ref len std 93.04 90.93 93.09
Bleu pred len mean 246.42 250.05 246.80
Bleu pred len std 92.99 90.89 92.96
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 25.11 27.48 25.18
Chrf score std 22.11 23.67 22.24
Ter score mean 89.07 86.14 88.79
Ter score std 26.62 28.31 26.49
Chrfpp score mean 25.11 27.48 25.18
Chrfpp score std 22.11 23.67 22.24
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.84 20.56 17.99
Bleu score std 18.05 19.71 16.23
Bleu 1gram precision mean 18.89 20.60 18.04
Bleu 1gram precision std 18.10 19.75 16.27
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 252.31 252.42 252.40
Bleu ref len std 91.64 91.62 91.68
Bleu pred len mean 255.15 255.25 255.12
Bleu pred len std 91.11 91.18 91.09
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 19.49 21.23 18.65
Chrf score std 18.37 20.02 16.58
Ter score mean 96.12 94.10 97.10
Ter score std 22.84 24.64 20.72
Chrfpp score mean 19.49 21.23 18.65
Chrfpp score std 18.37 20.02 16.58

### v3

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.28 15.08 14.55 15.03
Bleu score std 14.76 16.23 15.04 16.47
Bleu 1gram precision mean 14.34 15.15 14.63 15.09
Bleu 1gram precision std 14.82 16.29 15.11 16.53
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 177.82 179.16 178.20 178.23
Bleu ref len std 65.99 65.72 66.07 65.79
Bleu pred len mean 178.21 179.61 178.54 178.72
Bleu pred len std 66.06 65.78 66.13 65.84
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 14.77 15.59 15.05 15.53
Chrf score std 15.02 16.49 15.29 16.71
Ter score mean 102.45 101.53 102.09 101.54
Ter score std 18.88 20.35 18.95 20.63
Chrfpp score mean 14.77 15.59 15.05 15.53
Chrfpp score std 15.02 16.49 15.29 16.71
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 11.47 11.83 11.57 11.96
Bleu score std 10.75 11.92 11.66 11.66
Bleu 1gram precision mean 11.50 11.86 11.61 11.99
Bleu 1gram precision std 10.78 11.94 11.69 11.69
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 183.01 183.18 183.27 182.99
Bleu ref len std 66.11 66.21 66.10 66.26
Bleu pred len mean 185.58 185.77 185.77 185.58
Bleu pred len std 66.51 66.57 66.44 66.66
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05 0.05
Chrf score mean 11.95 12.32 12.05 12.45
Chrf score std 11.00 12.15 11.88 11.92
Ter score mean 106.35 106.05 106.22 105.88
Ter score std 15.12 16.63 15.96 16.35
Chrfpp score mean 11.95 12.32 12.05 12.45
Chrfpp score std 11.00 12.15 11.88 11.92
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 10.24 12.43 12.31 13.34
Bleu score std 11.68 14.39 13.71 15.66
Bleu 1gram precision mean 10.29 12.51 12.38 13.40
Bleu 1gram precision std 11.72 14.45 13.76 15.68
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 75.92 80.18 79.39 80.25
Bleu ref len std 31.48 32.63 32.60 32.88
Bleu pred len mean 79.72 84.07 83.38 84.51
Bleu pred len std 34.40 35.19 35.41 35.68
Bleu bp mean 0.98 0.99 0.99 0.99
Bleu bp std 0.11 0.08 0.09 0.08
Bleu ratio mean 1.04 1.05 1.05 1.05
Bleu ratio std 0.14 0.12 0.13 0.12
Chrf score mean 10.82 13.08 12.98 14.03
Chrf score std 12.05 14.82 14.16 16.07
Ter score mean 101.05 98.60 98.46 97.59
Ter score std 19.34 21.01 19.97 22.29
Chrfpp score mean 10.82 13.08 12.98 14.03
Chrfpp score std 12.05 14.82 14.16 16.07
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 13.67 14.77 14.40 15.66
Bleu score std 15.29 15.79 16.04 17.12
Bleu 1gram precision mean 13.75 14.86 14.48 15.73
Bleu 1gram precision std 15.31 15.84 16.08 17.15
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 75.67 81.10 78.32 79.15
Bleu ref len std 34.88 34.04 33.85 33.96
Bleu pred len mean 78.80 84.24 81.61 82.81
Bleu pred len std 36.63 35.88 35.88 36.02
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.05 0.04 0.04 0.03
Bleu ratio mean 1.04 1.04 1.04 1.05
Bleu ratio std 0.08 0.07 0.07 0.07
Chrf score mean 14.34 15.42 15.06 16.39
Chrf score std 15.65 16.10 16.38 17.44
Ter score mean 97.26 95.50 96.19 94.64
Ter score std 22.32 21.46 22.40 22.82
Chrfpp score mean 14.34 15.42 15.06 16.39
Chrfpp score std 15.65 16.10 16.38 17.44
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.21 18.44 18.60 18.55
Bleu score std 18.76 19.49 19.25 19.20
Bleu 1gram precision mean 17.35 18.54 18.70 18.65
Bleu 1gram precision std 18.95 19.62 19.38 19.32
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 83.21 85.58 85.50 85.22
Bleu ref len std 36.03 35.24 35.33 35.18
Bleu pred len mean 86.68 88.99 88.96 88.76
Bleu pred len std 37.81 37.00 37.10 36.95
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06 0.06
Chrf score mean 17.93 19.16 19.32 19.29
Chrf score std 19.10 19.82 19.56 19.56
Ter score mean 93.01 90.78 90.55 90.77
Ter score std 24.89 24.17 24.00 24.03
Chrfpp score mean 17.93 19.16 19.32 19.29
Chrfpp score std 19.10 19.82 19.56 19.56
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.50 14.90 14.70 15.21
Bleu score std 15.36 16.07 15.52 17.08
Bleu 1gram precision mean 14.58 14.98 14.78 15.30
Bleu 1gram precision std 15.49 16.20 15.65 17.21
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 84.04 84.57 84.06 84.32
Bleu ref len std 35.54 35.28 35.58 35.65
Bleu pred len mean 87.66 88.19 87.69 87.93
Bleu pred len std 37.74 37.49 37.81 37.86
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06 0.06
Chrf score mean 15.12 15.52 15.32 15.83
Chrf score std 15.65 16.35 15.81 17.32
Ter score mean 95.18 94.49 94.76 94.21
Ter score std 20.33 20.54 20.07 21.67
Chrfpp score mean 15.12 15.52 15.32 15.83
Chrfpp score std 15.65 16.35 15.81 17.32
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 12.04 12.98 12.21 12.89
Bleu score std 13.46 14.42 13.42 14.61
Bleu 1gram precision mean 12.08 13.03 12.24 12.94
Bleu 1gram precision std 13.51 14.47 13.46 14.67
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 86.46 86.23 86.48 86.25
Bleu ref len std 36.21 36.19 36.24 36.16
Bleu pred len mean 92.34 91.99 92.37 91.94
Bleu pred len std 39.90 39.68 39.94 39.65
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.02 0.01 0.02
Bleu ratio mean 1.07 1.07 1.07 1.07
Bleu ratio std 0.08 0.09 0.08 0.09
Chrf score mean 12.76 13.74 12.94 13.63
Chrf score std 13.86 14.81 13.83 14.97
Ter score mean 99.34 98.29 99.13 98.39
Ter score std 20.49 21.52 20.54 21.82
Chrfpp score mean 12.76 13.74 12.94 13.63
Chrfpp score std 13.86 14.81 13.83 14.97
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 22.68 25.41 23.57 26.08
Bleu score std 20.79 23.06 21.92 23.71
Bleu 1gram precision mean 22.75 25.47 23.64 26.13
Bleu 1gram precision std 20.83 23.10 21.96 23.74
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 203.28 234.13 207.02 214.82
Bleu ref len std 84.53 89.52 84.94 86.44
Bleu pred len mean 204.82 235.82 208.57 216.33
Bleu pred len std 85.31 90.00 85.73 86.78
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 23.33 26.10 24.25 26.76
Chrf score std 21.07 23.29 22.15 23.93
Ter score mean 91.40 88.33 90.51 87.39
Ter score std 25.52 27.99 26.91 28.75
Chrfpp score mean 23.33 26.10 24.25 26.76
Chrfpp score std 21.07 23.29 22.15 23.93
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 24.34 29.55 25.83 29.69
Bleu score std 20.94 24.54 21.80 24.75
Bleu 1gram precision mean 24.42 29.60 25.90 29.75
Bleu 1gram precision std 20.98 24.57 21.84 24.78
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 237.17 246.51 243.30 244.89
Bleu ref len std 90.05 87.83 87.19 87.45
Bleu pred len mean 238.76 248.14 244.78 246.06
Bleu pred len std 91.05 88.54 87.84 87.84
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.03 0.02 0.02 0.02
Bleu ratio mean 1.00 1.01 1.01 1.00
Bleu ratio std 0.03 0.02 0.03 0.02
Chrf score mean 25.05 30.30 26.55 30.42
Chrf score std 21.09 24.71 22.03 24.96
Ter score mean 89.75 83.28 87.62 83.14
Ter score std 26.10 29.48 26.32 29.68
Chrfpp score mean 25.05 30.30 26.55 30.42
Chrfpp score std 21.09 24.71 22.03 24.96
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 24.45 28.83 25.73 27.84
Bleu score std 22.85 26.47 23.16 25.72
Bleu 1gram precision mean 24.51 28.89 25.79 27.90
Bleu 1gram precision std 22.88 26.50 23.19 25.75
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 245.51 252.04 251.56 251.70
Bleu ref len std 96.46 91.08 91.27 91.21
Bleu pred len mean 245.92 252.59 251.95 252.32
Bleu pred len std 96.66 91.19 91.33 91.25
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.03 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02 0.02
Chrf score mean 25.04 29.44 26.35 28.48
Chrf score std 22.96 26.59 23.33 25.89
Ter score mean 89.16 83.80 87.36 84.94
Ter score std 27.90 31.46 27.52 30.57
Chrfpp score mean 25.04 29.44 26.35 28.48
Chrfpp score std 22.96 26.59 23.33 25.89
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 24.49 26.86 24.57 26.86
Bleu score std 21.92 23.53 22.03 23.85
Bleu 1gram precision mean 24.58 26.93 24.65 26.94
Bleu 1gram precision std 21.96 23.57 22.07 23.89
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 246.19 249.56 246.63 248.87
Bleu ref len std 93.04 90.93 93.09 90.42
Bleu pred len mean 246.42 250.05 246.80 249.24
Bleu pred len std 92.99 90.89 92.96 90.11
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 25.11 27.48 25.18 27.50
Chrf score std 22.11 23.67 22.24 24.04
Ter score mean 89.07 86.14 88.79 86.08
Ter score std 26.62 28.31 26.49 28.70
Chrfpp score mean 25.11 27.48 25.18 27.50
Chrfpp score std 22.11 23.67 22.24 24.04
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 18.84 20.56 17.99 21.03
Bleu score std 18.05 19.71 16.23 20.59
Bleu 1gram precision mean 18.89 20.60 18.04 21.07
Bleu 1gram precision std 18.10 19.75 16.27 20.63
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 252.31 252.42 252.40 252.39
Bleu ref len std 91.64 91.62 91.68 91.84
Bleu pred len mean 255.15 255.25 255.12 255.18
Bleu pred len std 91.11 91.18 91.09 91.29
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05 0.05
Chrf score mean 19.49 21.23 18.65 21.71
Chrf score std 18.37 20.02 16.58 20.89
Ter score mean 96.12 94.10 97.10 93.63
Ter score std 22.84 24.64 20.72 25.78
Chrfpp score mean 19.49 21.23 18.65 21.71
Chrfpp score std 18.37 20.02 16.58 20.89